OpenStudy (anonymous):

How would I multiply the following rational expression? ((m^2-1)/(m^2+1))*((m^4+2m^2+1)/(m-1))

OpenStudy (anonymous):

\[((m ^{2}-1)/(m ^{2}+1)) *((m ^{4}+2m ^{2}+1)/(m-1))\]

jhonyy9 (jhonyy9):

so there are m^2 -1 m^4 +2m^2 +1 -------- * ---------------- = m^2 +1 m-1 correct ?

OpenStudy (anonymous):

yup!

jhonyy9 (jhonyy9):

so how you see what is inside first parantheses ?

OpenStudy (anonymous):

(m-1)(m+1)

jhonyy9 (jhonyy9):

yes sure

jhonyy9 (jhonyy9):

and what is the second numerator

jhonyy9 (jhonyy9):

?

OpenStudy (anonymous):

(m+1)(m+1)

jhonyy9 (jhonyy9):

this is right too

OpenStudy (anonymous):

next bracket is (m-1)^2(m+1)^2/m-1

jhonyy9 (jhonyy9):

sorry but this how you have wrote i dont understand it

OpenStudy (anonymous):

((m-1)^2(m+1)^2)/(m-1)

jhonyy9 (jhonyy9):

why you have got this ?

OpenStudy (anonymous):

I dont know. Is the answer to the question (m-1)^2 (m+1) ?

jhonyy9 (jhonyy9):

so you have solved it step by step right than just you need to unit these steps together and will get the right result

OpenStudy (anonymous):

is (m-1)^2 (m+1) the correct answer?

jhonyy9 (jhonyy9):

(m-1)(m+1) (m^2 +1)^2 ---------- * ------------ = m^2 +1 m-1 came on just you need to simplifie it and will get the right result

jhonyy9 (jhonyy9):

courage please this is very easy

OpenStudy (anonymous):